# DATransNet: Dynamic Attention Transformer Network for Infrared Small Target Detection

### v3

@@ Abstract
Infrared small target detection (ISTD) is widely used in civilian and military applications. However, ISTD encounters several challenges, including the tendency for small and dim targets to be obscured by complex backgrounds. To address this issue, we propose the Dynamic Attention Transformer Network (DATransNet), which aims to extract and preserve edge information of small targets. DATransNet employs the Dynamic Attention Transformer (DATrans), simulating central difference convolutions (CDC) to extract and integrate gradient features with deeper features. Furthermore, we propose a global feature extraction module (GFEM) that offers a comprehensive perspective to prevent the network from focusing solely on details while neglecting the background information. We compare the network with state-of-the-art (SOTA) approaches, and the results demonstrate that our method performs effectively. Our source code is available at https://github.com/greekinRoma/DATransNet.
@@ I Introduction
INFRARED small target detection (ISTD) is critically important for numerous space-based computer vision applications. Although ISTD has been extensively studied, it remains a challenge. The considerable imaging distances associated with space-based infrared cameras often result in small targets that lack detailed texture and shape information.
Currently, the ISTD methods can be categorized into model-driven and data-driven methods. Model-driven methods include three main approaches. 1) Filter-based methods, such as the Tophat [1], enhance target visibility in complex backgrounds while suppressing background noise. 2) Methods based on the human visual system (HVS), such as the local contrast measure (LCM)[2] and multi-patch contrast measure (MPCM)[3], leverage human visual perception to improve local information. 3) Low-rank matrix decomposition and reconstruction methods, such as infrared patch image (IPI)[4], and Rewight infrared patch tensor (RIPT) [5], decompose small targets and backgrounds using low-rank matrices. However, these techniques require manual parameter settings and often struggle to adapt to complex scenarios.
With the advancement of deep learning, data-driven approaches have achieved substantial progress in infrared small target detection (ISTD). For example, the Asymmetric Context Modulation (ACM) network [6] introduces an asymmetric feature fusion technique, an alternative to conventional skip connections in U-Net. Building on this, DNANet [7] implements a multilayer nested architecture that supports progressive and adaptive interactions among feature layers. Moreover, UIUNet [8] enhances the detection of local target contrasts by integrating multiple U-Net structures and using interactive cross-attention mechanisms for feature fusion. Additionally, GSTUnet [9] merges vision Transformer technology with CNNs in the encoder to efficiently extract edge features from small targets. Generally, these data-driven methods surpass traditional model-driven approaches.
Some challenges faced by data-driven methods, such as the substantial number of parameters and the potential loss of local feature information. Our motivation focus on enhancing the feature information from small targets.
Although Data-driven methods achieve good performance, they suffer from a substantial number of parameters and potential loss of local information. Therefore, we propose some methods to improve the ISTD network, and the new network is called Dynamic Attention Transformer Network (DATransNet). DATransNet incorporates two key innovation components to improve feature extraction. First, we propose the Dynamic Attention Transformer (DATrans), which utilizes Transformers to simulate the central difference convolution (CDC) and enable the network to retain the detailed information of the inputs in deep layers and give dynamic weights to differences result from differenct directions. Second, the global feature extraction module (GFEM) solves the problem of lacking global background perception, thereby improving the ability to obtain contextual information.
@@
Figure 1: The overall structure of DATransNet. The red represents the downsampling module and the blue corresponds to the upsampling module.
@@
    We propose the DATransNet, which utilizes the DATrans to simulate the CDC with dynamic weights, extracting the edge and point information.
@@
    We propose the GFEM to improve the CNN from only focusing on local details to integrating these details with global information.
@@
    We conduct comparative experiments on the IRSTD-1K and NUDT-SIRST datasets, and the results show that DATransNet outperforms existing methods.
@@ II Methods
The overall architecture of DATransNet is illustrated in Fig. 1. DATrans is based on U-Net and implements DATrans and GFEM to extract detailed information and contextual features.
@@ II-A Dynamic Attention Transformer (DATrans)
CDCs are defined in Eq. 1 and are employed to extract edge and gradient features. Here, $c_{inp}$ represents the number of input channels, $o_{ixy}$ denotes the input value at the $i$-th channel and position $(x,y)$, and $b_{ijxy}$ refers to the $j$-th surrounding value around $o_{ixy}$. $out_{\hat{c}xy}$ denotes the output of the $\hat{c}$-th channel at position $(x,y)$. $w_{\hat{c}ijxy}$ is the weight applied to the difference $(o_{ixy}-b_{ijxy})$, which contributes to determining $out_{\hat{c}xy}$. A significant limitation of CDCs is the use of fixed weights for different directions, meaning the importance of each direction remains constant across different images.
@@
We can transform Eq. 1 into a fusion of Eq. 2 and Eq. 3. The process described in Eq. 2 is illustrated in Fig. 2. As shown in Fig. 2, we use edge convolution to extract the difference values from the input $T\in\mathbb{R}^{c_{inp}\times w\times h}$ and reshape them to produce a local difference matrix $D\in\mathbb{R}^{8c_{inp}\times wh}$. In Eq. 3, we utilize the weight matrix $M_{w}\in\mathbb{R}^{c_{o}\times 8c_{inp}}$ and the local difference matrix $D$ to obtain the final result $Out$. Here, $w$ denotes the width of the input, $h$ represents the input height, and $c_{o}$ signifies the number of output channels.
$D=reshape(Conv_{3\times 3}(T))\$ (2)
$Out=M_{w}D$ (3)
Figure 2: $D^{n}$ is obtained by performing edge detection convolution with dilation ratio of n on $T$ and unfolding the output.
When $M_{w}$ is influenced by the input $T$, the CDC could exhibit enhanced robustness. Based on the analysis above, we introduce the DATrans to simulate CDC and extract the details using dynamic weights from low-level layers for ISTD. The structure of DATrans is shown in Fig. 3.
@@
In varied heads of the DATrans, we employ different unfolding methods based on edge detection convolutions in Fig. 2 with varying dilation and reshape operations to get the unfolding .
In the various heads of DATrans, we apply different edge detection convolutions from Fig. 2 to $T$, utilizing varying dilation ratios. The outputs from these convolutions are then reshaped into $D^{n}\in\mathbb{R}^{8c_{inp}\times wh}$ where $n$ is the dilation ratio. Additionally, the input $T$ is reshaped to form $O\in\mathbb{R}^{c_{inp}\times wh}$. The token $O$ serves as the query, and the local difference matrix $D^{n}\in\mathbb{R}^{8c_{inp}\times wh}$ functions as both the key and the value.
$Q^{n}$ $=W^{n}_{Q}O,K^{n}=W^{n}_{K}D^{n},V=W^{n}_{V}D^{n}$ (4)
where $W^{n}_{Q}\in\mathbb{R}^{\frac{c_{o}}{m}\times c_{inp}}$, $W^{n}_{K}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times 8c_{inp}}$, $W^{n}_{V}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times 8c_{inp}}$ are the reshaped weights of convolutions and $m$ denotes the number of heads. We utilize the $Q^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times wh}$ and $K^{n}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times wh}$ to produce the attention matrix $M^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times\frac{8c_{inp}}{m}}$. $V^{n}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times wh}$ denotes the value of the transformer. The result $Out^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times wh}$ is defined as follows:
@@
where $M^{n}_{mix}\in\mathbb{R}^{\frac{c_{o}}{m}\times 8c_{inp}}$ denotes the dynamic matrix derived from a attention matrix $M^{n}$ and a fixed convolution weight matrix $W_{V}^{n}$. According to Eq. 3, $Out^{n}$ is the equivalent result of the CDC with dynamic weights and a dilation ratio of n. $Norm$ is the Instance Normalization. At last, we concatenate the results in varied heads, reshape them and use $1\times 1$ convolution to fuse them as shown in the Equation below:
@@
Overall, the gradient information from features at different scales can be fully extracted, enhancing the model’s ability to adapt to complex backgrounds.
@@ II-B Global Feature Extraction Module (GFEM)
Background information is crucial alongside detailed features in ISTD. However, perceiving background relies on global features, which the DATrans struggles to handle effectively. To address this limitation and enhance global information perception, we propose the GFEM, as illustrated in Fig. 4.
In GFEM, we utilize the attention mechanism to facilitate the transmission of information across distant pixels. This module includes two key aspects to broaden perception: First, we employ the non-local attention mechanism [10] to compute the spatial attention, enabling connectivity between pixels over long distances. Second, we incorporate a squeeze-and-excitation block [11] to enhance channel-wise perception. Finally, we concatenate the outputs from the global spatial and channel attention modules and apply convolution to fuse them, enabling GFEM to capture comprehensive feature representations.
Subsequently, each feature map acquires both spatial and channel global receptive fields, which are vital for effective object detection.
@@ II-C Loss Function
The loss function utilized in our model training is the soft intersection over union (SoftIoU) loss, where the predictions of DATransNet are mapped into the range of 0 and 1, using the Sigmoid function as shown in Eq. 7.
$p_{i,j}=\frac{1}{1+e^{-v_{i,j}}}$ (7)
where $p_{i,j}$ denotes the Sigmoid output of $v_{i,j}$, the model’s prediction at pixel $(i,j)$. Moreover, the SoftIoU loss is defined as:
@@
In this equation, $g_{i,j}$ denotes the ground truth value at pixel $(i,j)$, typically 0 or 1.
Figure 5: Visual results from varied data-driven methods. Red, green, and yellow boxes label detected targets, missed targets, and false alarms. The Red bounding box are zoom in, facilitating to auxiliary our to evaluation detection results. Our DATransNet can achieve high probability detection rate, and achieve precision localization and segmentation of their shapes.
@@ III Experiment and Analysis
This section first describes the experimental details and the evaluation metrics. Following with a series of ablation studies to verify the proposed modules. Finally, a comprehensive comparative experiments with other methods on qualitative and quantitative results demonstrate that our approach outperforms existing state-of-the-art (SOTA) methods.
@@ III-A Dataset and Evaluation Metrics
The datasets used for evaluating the DATransNet are NUDT-SIRST[7] and IRSTD-1K[12]. The evaluation metrics include mean intersection over union ($mIoU$), F-measure ($F_{1}$), the probability of detection ($P_{d}$), and false alarm rate ($F_{a}$). Meanwhile, we plot the ROC curves of various methods to evaluate the robust of DATransNet.
@@ III-B Experimental Environment and Parameter Settings
Our model is implemented by PyTorch framework, then training on NVIDIA GeForce RTX 4080 GPUs. We employ the Adam optimizer, starting with an initial learning rate of $5\times 10^{-4}$. This learning rate is reduced to $5\times 10^{-5}$ at epoch 200 and further decrease to $5\times 10^{-6}$ We use a batch size of 4 for training and run the model for 400 epochs.
@@ III-C Ablation Study
III-C1 Module-wise Performance Gain
In this ablation study, we begin with the most primitive Baseline, Which includes only feature extraction and upsampling with in the U-Net framework. Then, We test the performance of the DATransNet w/o each module, the results are shown in Tab. I. The results illustrate the effectiveness of each module. The performance of our network improves progressively with the inclusion of DATrans and GFEM. DATrans, which focuses on local difference features, enhances the model’s performance more effectively than GFEM, which primarily relies on global perception. In ISTD, the capability to learn gradient information is more critical than access to global context.
@@ III-C2 Studies of Dilation Rate for DATrans
As mentioned in Section II-A, DATrans employs varying dilation rates across different detection heads. Tab. II demonstrates the efficacy of this strategy, as models with diverse dilation ratios outperform those with a uniform rate. Furthermore, networks with dilation ratios of 1 and 3 yield superior results compared to those with other dilation ratios. Therefore, the dilation ratio significantly impacts target detection, and we choose the ratios of 1 and 3 as the dilation ratios for our network.
@@
Non-local SE Block mIoU(%) $F_{1}(\%)$ $P_{d}(\%)$ $F_{a}(10^{-6})$
$\checkmark$ -- 94.69 97.27 98.73 2.57
-- $\checkmark$ 94.53 97.19 98.23 1.75
$\checkmark$ $\checkmark$ 94.93 97.39 99.04 2.00

### v4

@@ Abstract
Infrared small target detection (ISTD) is widely used in civilian and military applications. However, ISTD encounters several challenges, including the tendency for small and dim targets to be obscured by complex backgrounds. To address this issue, we propose the Dynamic Attention Transformer Network (DATransNet), which aims to extract and preserve detailed information vital for small targets. DATransNet employs the Dynamic Attention Transformer (DATrans), simulating central difference convolutions (CDC) to extract gradient features. Furthermore, we propose a global feature extraction module (GFEM) that offers a comprehensive perspective to prevent the network from focusing solely on details while neglecting the global information. We compare the network with state-of-the-art (SOTA) approaches and demonstrate that our method performs effectively. Our source code is available at https://github.com/greekinRoma/DATransNet.
@@ I Introduction
INFRARED small target detection (ISTD) is vital for various fields. Currently, ISTD methods can be categorized into model-driven and data-driven methods.
Model-driven methods include three main approaches. 1) Filter-based methods, such as the Tophat [1]. 2) Methods based on the human visual system (HVS), such as the local contrast measure (LCM)[2] and multi-patch contrast measure (MPCM)[3]. 3) Low-rank matrix decomposition and reconstruction methods, such as infrared patch image (IPI)[4], and reweighted infrared patch tensor (RIPT) [5].
With the advancement of deep learning, data-driven approaches have achieved substantial progress in ISTD. For example, the Asymmetric Context Modulation (ACM) network [6] introduces asymmetric feature fusion, an alternative to conventional skip connections in U-Net. The dense nested attention network (DNANet) [7] implements a multilayer nested architecture that supports progressive and adaptive interactions between feature layers. Moreover, UIUNet [8] enhances the detection of local target contrasts by integrating multiple U-Net structures and using interactive cross-attention mechanisms for feature fusion. Additionally, Gated-shaped TransUnet (GSTUnet) [9] merges Vision Transformer with CNNs in the encoder to learn both global and local information. Receptive-field and Direction-induced Attention Network (RDIAN) [10] utilizes different receptive fields and multi-direction-guided attention to enhance the features of targets. Attentional Local Contrast Network (ALCNet) [11] designs a bottom-up attention modulation to strengthen the small target characteristics. Attention-guided Pyramid Context Network (AGPCNet) [12] divides the image into several patches and computes both global and local associations. Yuan et al. [13] propose the Spatial-channel Cross Transformer Network (SCTransNet), which uses the Spatial-channel Cross Transformer Blocks (SCTBs) to improve the capacity of global information modelling. Generally, these data-driven methods surpass traditional model-driven approaches.
Although data-driven methods achieve excellent performance, they often struggle with limited ability to capture details and weak global perceptual capabilities. To address these limitations, we propose a novel framework called the Dynamic Attention Transformer Network (DATransNet). DATransNet incorporates two key modules. We first introduce the Dynamic Attention Transformer (DATrans), which is good at capturing details. Then, the Global Feature Extractor Module (GFEM) gives our model a global perspective over the whole image.
@@
Figure 1: The overall structure of DATransNet. The red represents the upsampling stage (UpStage), and the blue corresponds to the downsampling stage (Stage).
@@
    We propose DATransNet, which utilizes DATrans to extract detailed information by simulating the central difference convolution (CDC) with dynamic weights.
@@
    We introduce GFEM to incorporate global information into our network.
@@
    We conducted comparative experiments on the IRSTD-1K and NUDT-SIRST datasets, and the results show that DATransNet outperforms many existing methods.
Figure 2: The first column is the original image, and the following 8 columns are the results of the original image using different edge convolutions across various directions.
@@ II Methods
The overall architecture of DATransNet is illustrated in Fig. 1. DATransNet is based on U-Net and implements DATrans and GFEM to extract detailed information and contextual features, respectively.
@@ II-A Dynamic Attention Transformer (DATrans)
As shown in Fig. 2, the targets in ISTD are small and dim and do not have complex texture information. Therefore, the details in the image are vital for ISTD. An effective approach to extracting the details is to take advantage of the differences between local pixels and their surroundings, such as CDCs. CDCs are defined in Eq. 1.
@@
Here, $c_{inp}$ represents the number of input channels, $o_{ixy}$ denotes the input value at the $i$-th channel and position $(x,y)$, and $b_{ijxy}$ refers to the $j$-th surrounding value around $o_{ixy}$. $out_{\hat{c}xy}$ denotes the output of the $\hat{c}$-th channel at position $(x,y)$. $w_{\hat{c}ijxy}$ is the weight applied to the difference $b_{ijxy}-o_{ixy}$.
The importance of different edge-convolution results varies from image to image. As shown in Fig. 2, the target in the upper image of the last column is clearly distinct, while that in the bottom row appears blurry. So, the importance of difference results in various directions is varied for different images. The weights $w_{\hat{c}ijxy}$ of the CDCs should be dynamically adjusted in response to changes in the input images.
We divide Eq. 1 into a fusion of Eq. 2, Eq. 3, and Eq. 4.
Eq. 2 and Eq. 3 describe the process of extracting difference features $T^{n}\in\mathbb{R}^{8c_{inp}\times(wh)}$, as shown in Fig. 3. We apply edge convolutions with a dilation ratio of $n$ to get $D^{n}\in\mathbb{R}^{8c_{inp}\times w\times h}$ from the input image $I\in\mathbb{R}^{c_{inp}\times w\times h}$, as outlined in Eq. 2. The $Diff$ refers to performing eight edge convolutions on the input image. In Eq. 3, the $Flatten$ reshapes $D^{n}$ into $T^{n}$. In addition, $w$ represents the input width, $h$ denotes the input height, and $c_{inp}$ denotes the number of input channels. In the process, we convert the spatial differences between eight neighborhoods to the dimension of the channel.
Then we use the weight matrix $M_{w}\in\mathbb{R}^{c_{o}\times 8c_{inp}}$ to obtain the output $Out\in\mathbb{R}^{8c_{inp}\times(wh)}$, as shown in Eq. 4, where $Reshape$ refers to the transformation process transforming $M_{w}T^{n}$ into $Out^{n}\in R^{c_{o}\times w\times h}$ and $c_{o}$ is the number of output channels.
$D^{n}=Diff(I)$ (2)
$T^{n}=Flatten(D^{n})$ (3)
$Out^{n}=Reshape(M_{w}T^{n})$ (4)
Figure 3: $T^{n}$ is derived by applying edge detection convolutions with a dilation ratio of $n$ to the image $I$, followed by flattening the output.
When $M_{w}$ in Eq. 4 changes with the input image $I$, the CDCs may also vary accordingly. Based on the analysis, we introduce the DATrans to simulate CDCs with dynamic weights for different directions. The structure of DATrans is shown in Fig. 4.
To improve the detection capability for a variety of targets, we utilize varying dilation ratios across different heads in DATrans. For each head, the difference feature ($T^{n}$) extraction process follows the same approach as that described in Eq. 2 and Eq. 3. We treat each channel with 8 surrounding information as a token.
@@
Furthermore, the input $I$ is reshaped to form $O\in\mathbb{R}^{c_{inp}\times wh}$, as shown in Eq. 5. $O$ serves as the query token. The key and value tokens come from $T^{n}\in\mathbb{R}^{8c_{inp}\times wh}$.
$O=Flatten(I)$ (5)
$Q^{n}$ $=W^{n}_{Q}O,K^{n}=W^{n}_{K}T^{n},V=W^{n}_{V}T^{n}$ (6)
where $W^{n}_{Q}\in\mathbb{R}^{\frac{c_{o}}{m}\times c_{inp}}$, $W^{n}_{K}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times 8c_{inp}}$, $W^{n}_{V}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times 8c_{inp}}$ are the learnable weights and $m$ denotes the number of heads. We use $Q^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times wh}$ and $K^{n}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times wh}$ to produce the attention matrix $M^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times\frac{8c_{inp}}{m}}$. $V^{n}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times wh}$ is multiplied by $M^{n}$. The result $Out^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times wh}$ is defined as follows:
@@
where $M^{n}_{mix}\in\mathbb{R}^{\frac{c_{o}}{m}\times 8c_{inp}}$ denotes the dynamic matrix derived from an attention matrix $M^{n}$ and a learnable weight matrix $W_{V}^{n}$. $Norm$ refers to Instance Normalization, which normalizes the similarity matrix for each instance on the similarity maps, ensuring smooth gradient propagation. According to Eq. 4 and 7, $Out^{n}$ is the equivalent result of the CDC with dynamic weights.
At last, we concatenate the results of varied heads which have different dilation ratios, reshape them and use a $1\times 1$ convolution to fuse them to get the final output $Out$. In addition, $m$ is the number of heads.
@@ II-B Global Feature Extraction Module (GFEM)
Background information is crucial alongside the detailed features of small targets in ISTD. However, background information is based on a global perspective of the whole image. So, we propose the GFEM, as illustrated in Fig. 5.
@@
In GFEM, we incorporate the attention mechanism to provide our model with a broader perspective. This module includes three key steps: First, we utilize the non-local attention mechanism [14] to compute spatial attention over the deepest feature map, which is significantly smaller in size compared to the input image. Initially, we apply a $1\times 1$ convolution to reduce the number of channels from $c$ in $X\in\mathbb{R}^{c\times w\times h}$ to $c^{\prime}$ in $Q\in\mathbb{R}^{c^{\prime}\times w\times h}$, $K\in\mathbb{R}^{c^{\prime}\times w\times h}$, and $V\in\mathbb{R}^{c^{\prime}\times w\times h}$. Besides, $w$ and $h$ denote the width and height of the feature map input to the non-local module. Then, we reshape the feature maps and compute $Y\in\mathbb{R}^{c\times w\times h}$ according to the following equation:
$Y=Conv_{1\times 1}(Reshape(V^{\prime}\cdot Softmax(Q^{\prime}\cdot(K^{\prime})^{T})))$
where $Q^{\prime}\in\mathbb{R}^{c^{\prime}\times(wh)}$, $K^{\prime}\in\mathbb{R}^{c^{\prime}\times(wh)}$, and $V^{\prime}\in\mathbb{R}^{c^{\prime}\times(wh)}$ are the results of reshaping $Q$, $K$, and $V$. $Softmax$ is softmax layer, and $Conv_{1\times 1}$ is a convolution with a $1\times 1$ kernel, to increase the number of channels. Second, we incorporate a squeeze and excitation block [15] to enhance channel-wise perception. Finally, we concatenate the results from the global spatial and channel attention modules and employ convolution to fuse them, enabling GFEM to capture comprehensive feature representations.
Subsequently, the output feature map acquires both spatial and channel global receptive fields, which are vital for effective object detection.
@@ II-C Loss Function
The loss function utilized in our model training is the soft intersection over union (SoftIoU) loss, as shown in Eq. 9.
@@
where $g_{i,j}$ and $p_{i,j}$ denote the ground truth and the output of our network at the $(i,j)$, respectively.
Figure 6: Visual results from varied data-driven methods. The red, green and yellow boxes represent detected targets, missed targets, and false alarms.
@@ III Experiment and Analysis
This section describes the experimental details and the evaluation metrics, followed by a series of ablation studies to verify the proposed modules. Finally, a comprehensive comparative experiment with other methods on qualitative and quantitative results demonstrates that our approach outperforms existing state-of-the-art (SOTA) methods.
@@ III-A Dataset and Evaluation Metrics
The datasets used for evaluating the DATransNet are NUDT-SIRST[7] and IRSTD-1K[16]. The NUDT-SIRST dataset consists of 1327 images whose resolution is 256×256, with a split of 332 images for testing, 332 for validation, and 663 for training. The IRSTD-1K dataset contains 1001 images with a resolution of 512×512. We select 101 images for testing, 100 images for validation, and 800 images for training. The evaluation metrics include mean intersection over union ($mIoU$), F1-measure ($F_{1}$), the probability of detection ($P_{d}$), and false alarm rate ($F_{a}$).
@@ III-B Experimental Environment and Parameter Settings
All models are built on the PyTorch framework and are trained on an NVIDIA GeForce RTX 4080 GPU. We employ the Adam optimizer, starting with an initial learning rate of $5\times 10^{-4}$. This learning rate is reduced to $5\times 10^{-5}$ at epoch 200 and further decreases to $5\times 10^{-6}$. The batch size is 4, and there are 400 epochs.
@@ III-C Ablation Study
III-C1 Studies of Module-wise Performance Gain
In this ablation study, we begin with the baseline. Then, we test the performance of the DATransNet’s module, and the results are shown in Tab. I. The integration of DATrans and GFEM leads to a progressive enhancement in the performance of our network..
@@ III-C2 Studies of Dilation Rate for DATrans
As mentioned in Section II-A, DATrans employs varying dilation rates across different detection heads. Tab. II demonstrates the efficacy of this strategy, as models with diverse dilation ratios outperform those with a single rate. Furthermore, networks with dilation ratios of 1 and 3 yield superior results.
@@
Non-local SE Block Params(M) GFLOPs(G) mIoU(%) $F_{1}(\%)$
-- -- 3.70 10.82 94.25 97.03
$\checkmark$ -- 4.03 10.89 94.69 97.27
-- $\checkmark$ 4.02 10.89 94.53 97.19
$\checkmark$ $\checkmark$ 4.04 10.90 94.93 97.39
Model Metrics NUDT-SIRST IRSTD-1K
Params(M) mIoU(%) $F_{1}(\%)$ $P_{d}(\%)$ $F_{a}(10^{-6})$ mIoU(%) $F_{1}(\%)$ $P_{d}(\%)$ $F_{a}(10^{-6})$
ACM[6] 0.40 70.97 82.99 97.67 7.33 64.09 78.11 88.55 17.21
RDIAN[10] 0.90 87.78 93.47 97.88 9.67 65.68 79.30 91.24 10.31
ALCNet[11] 0.37 92.45 96.08 98.94 2.62 65.05 78.60 90.57 19.42
DNANet[7] 4.69 93.73 96.70 99.25 4.55 66.73 80.08 89.22 7.97
UIU-Net[8] 50.54 94.11 96.96 97.98 0.74 67.94 80.90 90.57 24.44
SCTransNet[13] 11.19 93.83 96.77 98.20 1.56 65.83 79.40 90.57 14.50
Ours 4.04 94.93 97.39 99.04 2.00 68.56 81.34 93.60 24.96